A 2200 kg truck encounters an average friction force of 785 N at interstate speeds. Suppose the truck accelerates from 25 m/s to 35 m/s over a distance of 350 m. Determine the amount of force generated by the truck’s drive-train in order to produce this result. 


1
Expert's answer
2021-12-03T12:43:59-0500

"a=\frac{v_2^2-v_1^2}{2s}=0.86~\frac m {s^2},"

"F'=ma=1900~N,"

"F=F'+F_0=2700~N."